OpenStudy (anonymous):

will fan and medal Which expressions are equal to the expression –2.7 – 16.4? Choose all answers that are correct. A. –2.7 – (–16.4) B. –16.4 + (–2.7) C. –16.4 – (–2.7) D. –2.7 + (–16.4)

OpenStudy (anonymous):

A and C

OpenStudy (anonymous):

right?

OpenStudy (igreen):

-2.7 - 16.4 = -19.1 –2.7 – (–16.4) = 13.7 –16.4 + (–2.7) = -19.1 –16.4 – (–2.7) = -13.7 –2.7 + (–16.4) = -19.1

OpenStudy (igreen):

No, it's B and D.

OpenStudy (anonymous):

oh.

OpenStudy (anonymous):

Yeah! it's B and D

OpenStudy (anonymous):

ohhhh. i see

OpenStudy (anonymous):

okay i will just fan all of you that helped

OpenStudy (anonymous):

and medal